OpenStudy (anonymous):

f(x) = 2/x^2-2x-3 Graph (1 point) Domain and Range (1 point): _____________________ x and y Intercept(s) (1 point): _____________________ Horizontal Asymptote(s) (1 point): ___________________ Vertical Asymptote(s) (1 point): ____________________

OpenStudy (anonymous):

put denominator = 0 and factorize it x^2-2x-3 = 0 x^2 -3x+x -3 = 0 x(x-3) +1(x-3) = 0 (x+1)(x-3) = 0 x = -1 and x = 3 thus the domain is all values of x except x = -1 and x = 3